OpenStudy (anonymous):

HELP :) MEDALS _THe perimeter of a rectangle is 54 cm. The area of the same rectangle is 176 cm2.What are the dimensions of the rectangle ? a.11 cm by 16 cm b.8cm by 22 cm c.5.5 cm by 32 cm d.4cm by 44 cm

OpenStudy (anonymous):

I have no clue how to solve this can you explain it to me ?

OpenStudy (anonymous):

i'll write it then send a pic

OpenStudy (anonymous):

Ok ty I've been trying to solve it for one hr ugg

OpenStudy (anonymous):

its okay. i was never a fan of word problems

OpenStudy (anonymous):

Ty your a angel :)

OpenStudy (anonymous):

OMG your amazing ty :) I have one more you think u can help ?

OpenStudy (anonymous):

possibly

OpenStudy (anonymous):

lay it on me. i'll give it my best

OpenStudy (anonymous):

Medal For Help :0) Which expression gives the solution of -5+2x2=6x? a. 2+ 4-(4)(6)(-5) =12 b.-5 + 25-(4)(2)(6) =-10 c. -6+ 36-(4)(2)(-5) =4 d. 6+ 36-(4)(2)(5) =4

OpenStudy (anonymous):

are you sure you gave me the right numbers?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

if the answers are in terms of the quadratic formula, and there is supposed to be a square root sign in the answers, then i may have an answer.....

OpenStudy (anonymous):

Yes I couldn't find the symbols

OpenStudy (anonymous):

\[\frac{ 6+\sqrt{36-(4)(2)(-5)} }{ 4 } = 4\]

OpenStudy (anonymous):

Yes your awesome , you should be a nath teacher :) Ty A million .If you every need help in any other subjects please do HOLLAR :)

OpenStudy (anonymous):

You're welcome. and thank you; I try. I will if i need help